# On second-order variational analysis of variational convexity of prox-regular functions

### v1

@@ 4 On neighborhood-based second-order characterizations of variational $s$-convexity
$\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)=\mathrm{gph}\,\partial\psi\cap(U\times V)\quad\mbox{and $f(x)=\psi(x)$ at the common elements $(x,x^{*})$}.$
@@
$\mathrm{gph}_{\widehat{\rho}}\,\partial\psi\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}.$
@@
$\mathrm{gph}_{\tilde{\rho}}\,\partial\psi\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}$
and, by possibly shrinking $\tilde{U}\times\tilde{V}$, we may assume that $\tilde{U}\times\tilde{V}\subset U\times V$. Then for any $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})$ we have $(x,x^{*})\in\mathrm{gph}_{\tilde{\rho}}\,\partial\psi\cap(U\times V)$ implying $\widehat{\psi}(x)=\psi(x)=f(x)<\tilde{\rho}$ and $\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})$. Since $\psi\leq f$ on $\tilde{U}\subset U$, we have $L_{\psi}(\tilde{\rho})\cap\tilde{U}\supset L_{f}(\tilde{\rho})\cap\tilde{U}$ and, together with $L_{f}(\rho)\supset L_{f}(\tilde{\rho})$, we obtain that
@@
$=\mathrm{gph}\,\partial f\cap(L_{f}(\rho)\times\mathbb{R}^{n})\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})=\mathrm{gph}_{\tilde{\rho}}\,\partial f\cap(\tilde{U}\times\tilde{V})$
@@
Let $y^{*}\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$ be arbitrarily fixed and note that $\partial\psi(x)=\partial f(x)+y^{*}+t(x-\bar{x})$. Next we choose some sufficiently small neighborhood $\widehat{U}\times\widehat{V}$ of $(\bar{x},{\bar{x}^{\ast}}+y^{*})$ such that $\widehat{U}\subset U$, $\widehat{V}-y^{*}-t(U-\bar{x})\subset V$ and $\langle y^{*},x-\bar{x}\rangle+\frac{t}{2}\|x-\bar{x}\|^{2}>-\epsilon$, $x\in\widehat{U}$, where $\epsilon:=(\rho-f(\bar{x}))/2$. Since $\widehat{f}\leq f$ on $U$, it follows that $\widehat{\psi}\leq\psi$ on $U\supset\widehat{U}$. Further, since $\widehat{f}-\frac{s}{2}\|\cdot\|^{2}$ is convex on $U$, $\widehat{\psi}-\frac{s+t}{2}\|\cdot\|^{2}$ is also convex there and we can conclude that $\widehat{\psi}$ is subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}+y^{*}$. Hence, by possibly shrinking $\widehat{U}\times\widehat{V}$, we may assume that $\widehat{\psi}(x)<\widehat{\rho}:=f(\bar{x})+\epsilon$ whenever $(x,x^{*})\in\widehat{U}\times\widehat{V}$. Then for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})$ we have $(x,x^{*}-y^{*}-t(x-\bar{x}))\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ and therefore $(x,x^{*}-y^{*}-t(x-\bar{x}))\in\mathrm{gph}\,\partial f$ and $f(x)=\widehat{f}(x)$. Hence $(x,x^{*})\in\mathrm{gph}\,\partial\psi(x)\cap(\widehat{U}\times\widehat{V})$ and $\psi(x)=\widehat{\psi}(x)<\widehat{\rho}$. On the other hand, if $(x,x^{*})\in\mathrm{gph}_{\widehat{\rho}}\,\partial\psi(x)\cap(\widehat{U}\times\widehat{V})$ we obtain that
@@
$\mathrm{gph}_{\widehat{\rho}}\,\partial\psi\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}$
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad z^{*}\in\widehat{D}_{f}^{*}(\partial f)(x,x^{*})(z),\ (x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$ (4.10)
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad z^{*}\in D_{f}^{*}(\partial f)(x,x^{*})(z),\ (x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$ (4.11)
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad(z,z^{*})\in L\in{\mathcal{S}}_{f}(\partial f)(x,x^{*}),\ (x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$ (4.13)
@@
    whenever $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\cap V)$.
@@
$\widehat{f}\leq f\mbox{ on }\widehat{U},\ \mathrm{gph}_{\widehat{\rho}}\,\partial f\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\widehat{U}\times\widehat{V})\mbox{ and $f(x)=\widehat{f}(x)$ at the common elemnts $(x,x^{*})$.}$ (4.14)
As pointed out in Remark 2.8(i), variational convexity of $f$ implies prox-regularity with parameter $r=0$ and some $\epsilon>0$. By possibly shrinking $\epsilon$ we may assume that ${\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset\widehat{U}\times\widehat{V}$ and $\rho:=f(\bar{x})+\epsilon\leq\widehat{\rho}$. The function $\widehat{f}$ is convex on $\widehat{U}$ and consequently variationally convex, prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$. Thus we may use the characterization [8, Theorem 5.2] of variational convexity to obtain the existence of a neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that
@@
where $\Phi(x,x^{*})$ is either $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})$ or $D^{*}(\partial\widehat{f})(x,x^{*})$. By possibly shrinking the neighborhood $U\times V$, we may assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})$ and $\widehat{f}(x)<\rho$ whenever $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$. Thus for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ we have $(x,x^{*})\in\mathrm{gph}\,\partial f$ and $f(x)=\widehat{f}(x)<f(\bar{x})+\epsilon$ implying $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\subset\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$. On the other hand, for every $(x,x^{*})\in\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)\subset\mathrm{gph}_{\widehat{\rho}}\,\partial f\cap(\widehat{U}\times\widehat{V})$ we have $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ and $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$ follows. Hence $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})=\widehat{D}^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ and $D^{*}(\partial\widehat{f})(x,x^{*})=D^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})$, $(x,x^{*})\in\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$ and the statements (ii) and (iii) follow from (4.15) and Proposition 3.5.
@@
Next assume that (ii) or (iii) hold with $s=0$. Since $f$ is assumed to be prox-regular at $\bar{x}$ for ${\bar{x}^{\ast}}$ with some parameter values $r\geq 0$ and $\epsilon>0$, by taking into account Remark 2.8(i), $f$ is also variationally $(-r)$-convex there and hence there exists an lsc function $\widehat{f}$ together with some open convex $\widehat{U}\times\widehat{V}$ of $(\bar{x},{\bar{x}^{\ast}})$ and some $\widehat{\rho}>f(\bar{x})$ such that $\widehat{f}+\frac{r}{2}\|\cdot\|^{2}$ is convex on $\widehat{U}$ and (4.14) holds. By possibly shrinking the neighborhood $U\times V$ in (ii) or (iii) and lowering $\rho$ and $\epsilon$, by utilizing subdifferential continuity of $\widehat{f}$ at $\bar{x}$ for ${\bar{x}^{\ast}}$ we can assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset\widehat{U}\times\widehat{V}$ and $\widehat{f}(x)<\rho\leq f(\bar{x})+\epsilon\leq\widehat{\rho}$ whenever $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$. Using the same arguments as above, we conclude that $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)=\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)$ and $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})=\widehat{D}^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})=\widehat{D}_{f}^{*}(\partial f)(x,x^{*})$ and $D^{*}(\partial\widehat{f})(x,x^{*})=D^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})=D_{f}^{*}(\partial f)(x,x^{*})$ for all $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)$ follows. Hence (4.15) is valid and, since $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$, by [8, Theorem 5.2] we obtain that $\widehat{f}$ is variationally convex at $\bar{x}$ for ${\bar{x}^{\ast}}$. Now (i) follows from Lemma 4.1.
@@
If (i) holds then $\partial f$ is $f$-locally monotone around $(\bar{x},{\bar{x}^{\ast}})$ and hence $\mathcal{T}^{f}_{\epsilon}$ is monotone around $(\bar{x},{\bar{x}^{\ast}})$ for all sufficiently small $\epsilon>0$ and by [13, Corollary 5.8] this is equivalent with the existence of some neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that for all $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ the graphical derivative $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is a monotone mapping. In order to apply [13, Corollary 5.8] note that at points $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ the mapping $\mathcal{T}^{f}_{\epsilon}$ is in fact proto-differentiable because of [20, Proposition 13.46] and [3, Remark 3.18]. From Proposition 3.5 we conclude that $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})\in{\cal Z}_{nn}$ for every $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ and thus $(x,x^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(f(\bar{x})+\epsilon))\times V)$. Further, monotonicity of $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ implies $\langle z^{*},z\rangle\geq 0$, $(z,z^{*})\in\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})$. This proves the implication (i)$\Rightarrow$(iv) with $\rho=f(\bar{x})+\epsilon$.
@@
Conversely, if (iv) holds then we may choose the parameter $\epsilon$ for prox-regularity so small that ${\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset U\times V$ and $f(\bar{x})+\epsilon<\rho$. Hence, $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ implies $(x,x^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(\rho))\times V)$ and $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})\in{\cal Z}_{nn}$ by virtue of Proposition 3.5. Since $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is a subspace, it follows from (4.12) that $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is monotone for these $(x,x^{*})$ and consequently $\mathcal{T}^{f}_{\epsilon}$ is locally monotone at $(\bar{x},{\bar{x}^{\ast}})$ by [13, Corollary 5.8] which implies that $\partial f$ has an $f$-attentive monotone localization around $(\bar{x},{\bar{x}^{\ast}})$ and consequently $f$ is variationally convex at $\bar{x}$ for ${\bar{x}^{\ast}}$.

### v2

@@ 4 On neighborhood-based second-order characterizations of variational $s$-convexity
$\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)=\mathrm{gph}\,\partial\psi\cap(U\times V)\quad\mbox{and $f(x)=\psi(x)$ at the common elements $(x,x^{*})$}.$
@@
$\mathrm{gph}\,\partial_{\widehat{\rho}}\psi\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}.$
@@
$\mathrm{gph}\,\partial_{\tilde{\rho}}\psi\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}$
and, by possibly shrinking $\tilde{U}\times\tilde{V}$, we may assume that $\tilde{U}\times\tilde{V}\subset U\times V$. Then for any $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})$ we have $(x,x^{*})\in\mathrm{gph}\,\partial_{\tilde{\rho}}\psi\cap(U\times V)$ implying $\widehat{\psi}(x)=\psi(x)=f(x)<\tilde{\rho}$ and $\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})$. Since $\psi\leq f$ on $\tilde{U}\subset U$, we have $L_{\psi}(\tilde{\rho})\cap\tilde{U}\supset L_{f}(\tilde{\rho})\cap\tilde{U}$ and, together with $L_{f}(\rho)\supset L_{f}(\tilde{\rho})$, we obtain that
@@
$=\mathrm{gph}\,\partial f\cap(L_{f}(\rho)\times\mathbb{R}^{n})\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})=\mathrm{gph}\,\partial_{\tilde{\rho}}f\cap(\tilde{U}\times\tilde{V})$
@@
Let $y^{*}\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$ be arbitrarily fixed and note that $\partial\psi(x)=\partial f(x)+y^{*}+t(x-\bar{x})$. Next we choose some sufficiently small neighborhood $\widehat{U}\times\widehat{V}$ of $(\bar{x},{\bar{x}^{\ast}}+y^{*})$ such that $\widehat{U}\subset U$, $\widehat{V}-y^{*}-t(U-\bar{x})\subset V$ and $\langle y^{*},x-\bar{x}\rangle+\frac{t}{2}\|x-\bar{x}\|^{2}>-\epsilon$, $x\in\widehat{U}$, where $\epsilon:=(\rho-f(\bar{x}))/2$. Since $\widehat{f}\leq f$ on $U$, it follows that $\widehat{\psi}\leq\psi$ on $U\supset\widehat{U}$. Further, since $\widehat{f}-\frac{s}{2}\|\cdot\|^{2}$ is convex on $U$, $\widehat{\psi}-\frac{s+t}{2}\|\cdot\|^{2}$ is also convex there and we can conclude that $\widehat{\psi}$ is subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}+y^{*}$. Hence, by possibly shrinking $\widehat{U}\times\widehat{V}$, we may assume that $\widehat{\psi}(x)<\widehat{\rho}:=f(\bar{x})+\epsilon$ whenever $(x,x^{*})\in\widehat{U}\times\widehat{V}$. Then for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})$ we have $(x,x^{*}-y^{*}-t(x-\bar{x}))\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ and therefore $(x,x^{*}-y^{*}-t(x-\bar{x}))\in\mathrm{gph}\,\partial f$ and $f(x)=\widehat{f}(x)$. Hence $(x,x^{*})\in\mathrm{gph}\,\partial\psi(x)\cap(\widehat{U}\times\widehat{V})$ and $\psi(x)=\widehat{\psi}(x)<\widehat{\rho}$. On the other hand, if $(x,x^{*})\in\mathrm{gph}\,\partial_{\widehat{\rho}}\psi(x)\cap(\widehat{U}\times\widehat{V})$ we obtain that
@@
$\mathrm{gph}\,\partial_{\widehat{\rho}}\psi\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{\psi}\cap(\widehat{U}\times\widehat{V})\quad\mbox{and $\psi(x)=\widehat{\psi}(x)$ at the common elements $(x,x^{*})$}$
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad z^{*}\in\widehat{D}_{f}^{*}(\partial f)(x,x^{*})(z),\ (x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$ (4.10)
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad z^{*}\in D_{f}^{*}(\partial f)(x,x^{*})(z),\ (x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$ (4.11)
@@
    $\langle z^{*},z\rangle\geq s\|z\|^{2}\quad\mbox{ whenever }\quad(z,z^{*})\in L\in{\mathcal{S}}_{f}(\partial f)(x,x^{*}),\ (x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$ (4.13)
@@
    whenever $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\cap V)$.
@@
$\widehat{f}\leq f\mbox{ on }\widehat{U},\ \mathrm{gph}\,\partial_{\widehat{\rho}}f\cap(\widehat{U}\times\widehat{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\widehat{U}\times\widehat{V})\mbox{ and $f(x)=\widehat{f}(x)$ at the common elemnts $(x,x^{*})$.}$ (4.14)
As pointed out in Remark 2.8(i), variational convexity of $f$ implies prox-regularity with parameter $r=0$ and some $\epsilon>0$. By possibly shrinking $\epsilon$ we may assume that ${\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset\widehat{U}\times\widehat{V}$ and $\rho:=f(\bar{x})+\epsilon\leq\widehat{\rho}$. The function $\widehat{f}$ is convex on $\widehat{U}$ and consequently variationally convex, prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$. Thus we may use the characterization [9, Theorem 5.2] of variational convexity to obtain the existence of a neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that
@@
where $\Phi(x,x^{*})$ is either $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})$ or $D^{*}(\partial\widehat{f})(x,x^{*})$. By possibly shrinking the neighborhood $U\times V$, we may assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})$ and $\widehat{f}(x)<\rho$ whenever $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$. Thus for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ we have $(x,x^{*})\in\mathrm{gph}\,\partial f$ and $f(x)=\widehat{f}(x)<f(\bar{x})+\epsilon$ implying $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\subset\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$. On the other hand, for every $(x,x^{*})\in\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)\subset\mathrm{gph}\,\partial_{\widehat{\rho}}f\cap(\widehat{U}\times\widehat{V})$ we have $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ and $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$ follows. Hence $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})=\widehat{D}^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ and $D^{*}(\partial\widehat{f})(x,x^{*})=D^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})$, $(x,x^{*})\in\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$ and the statements (ii) and (iii) follow from (4.15) and Proposition 3.5.
@@
Next assume that (ii) or (iii) hold with $s=0$. Since $f$ is assumed to be prox-regular at $\bar{x}$ for ${\bar{x}^{\ast}}$ with some parameter values $r\geq 0$ and $\epsilon>0$, by taking into account Remark 2.8(i), $f$ is also variationally $(-r)$-convex there and hence there exists an lsc function $\widehat{f}$ together with some open convex $\widehat{U}\times\widehat{V}$ of $(\bar{x},{\bar{x}^{\ast}})$ and some $\widehat{\rho}>f(\bar{x})$ such that $\widehat{f}+\frac{r}{2}\|\cdot\|^{2}$ is convex on $\widehat{U}$ and (4.14) holds. By possibly shrinking the neighborhood $U\times V$ in (ii) or (iii) and lowering $\rho$ and $\epsilon$, by utilizing subdifferential continuity of $\widehat{f}$ at $\bar{x}$ for ${\bar{x}^{\ast}}$ we can assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset\widehat{U}\times\widehat{V}$ and $\widehat{f}(x)<\rho\leq f(\bar{x})+\epsilon\leq\widehat{\rho}$ whenever $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$. Using the same arguments as above, we conclude that $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)=\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)$ and $\widehat{D}^{*}(\partial\widehat{f})(x,x^{*})=\widehat{D}^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})=\widehat{D}_{f}^{*}(\partial f)(x,x^{*})$ and $D^{*}(\partial\widehat{f})(x,x^{*})=D^{*}\mathcal{T}^{f}_{\epsilon}(x,x^{*})=D_{f}^{*}(\partial f)(x,x^{*})$ for all $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)$ follows. Hence (4.15) is valid and, since $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$, by [9, Theorem 5.2] we obtain that $\widehat{f}$ is variationally convex at $\bar{x}$ for ${\bar{x}^{\ast}}$. Now (i) follows from Lemma 4.1.
@@
If (i) holds then $\partial f$ is $f$-locally monotone around $(\bar{x},{\bar{x}^{\ast}})$ and hence $\mathcal{T}^{f}_{\epsilon}$ is monotone around $(\bar{x},{\bar{x}^{\ast}})$ for all sufficiently small $\epsilon>0$ and by [15, Corollary 5.8] this is equivalent with the existence of some neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that for all $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ the graphical derivative $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is a monotone mapping. In order to apply [15, Corollary 5.8] note that at points $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ the mapping $\mathcal{T}^{f}_{\epsilon}$ is in fact proto-differentiable because of [22, Proposition 13.46] and [4, Remark 3.18]. From Proposition 3.5 we conclude that $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})\in{\cal Z}_{nn}$ for every $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ and thus $(x,x^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(f(\bar{x})+\epsilon))\times V)$. Further, monotonicity of $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ implies $\langle z^{*},z\rangle\geq 0$, $(z,z^{*})\in\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})$. This proves the implication (i)$\Rightarrow$(iv) with $\rho=f(\bar{x})+\epsilon$.
@@
Conversely, if (iv) holds then we may choose the parameter $\epsilon$ for prox-regularity so small that ${\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})\subset U\times V$ and $f(\bar{x})+\epsilon<\rho$. Hence, $(x,x^{*})\in{\cal O}_{\mathcal{T}^{f}_{\epsilon}}\cap(U\times V)$ implies $(x,x^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(\rho))\times V)$ and $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})=T_{\mathrm{gph}\,\partial f}^{f}(x,x^{*})\in{\cal Z}_{nn}$ by virtue of Proposition 3.5. Since $\mathrm{gph}\,D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is a subspace, it follows from (4.12) that $D\mathcal{T}^{f}_{\epsilon}(x,x^{*})$ is monotone for these $(x,x^{*})$ and consequently $\mathcal{T}^{f}_{\epsilon}$ is locally monotone at $(\bar{x},{\bar{x}^{\ast}})$ by [15, Corollary 5.8] which implies that $\partial f$ has an $f$-attentive monotone localization around $(\bar{x},{\bar{x}^{\ast}})$ and consequently $f$ is variationally convex at $\bar{x}$ for ${\bar{x}^{\ast}}$.